OpenStudy (yacoub1993):

A cheetah ran 300 feet in 2.92 seconds. What was the cheetah’s average speed in miles per hour?

OpenStudy (yttrium):

How many seconds are there in one hr? How many feet are there in one mile?

OpenStudy (yacoub1993):

60 sec

OpenStudy (yacoub1993):

5280 feet in one mile

OpenStudy (yttrium):

are you sure with 60 sec?

OpenStudy (yacoub1993):

60 min i mean sory

OpenStudy (yttrium):

Then you can apply the formula S = distance/time to find for speed, then convert it. Got?

OpenStudy (yacoub1993):

let me try it

OpenStudy (yttrium):

sure :)

OpenStudy (yacoub1993):

so the distance is 300/5280 =0.5681818182

OpenStudy (yacoub1993):

and what is suppose to be the time 60*60 =0.0243333

OpenStudy (yacoub1993):

i am not sure with the time

OpenStudy (yttrium):

Yes, the dist is 300/5280 and the time is 2.92/3600

OpenStudy (yttrium):

Is this what you did?

OpenStudy (yacoub1993):

why 3600

OpenStudy (yacoub1993):

ohh got it wait let me try to solve it

OpenStudy (yacoub1993):

yes i tried that way distance / time = 0.05681818182 / 10512 = 5.405079

OpenStudy (yacoub1993):

there is something going wrong

OpenStudy (yacoub1993):

@Yttrium could you explain a little more clear Please

OpenStudy (anonymous):

A cheetah ran 300 feet in 2.92 seconds. What is the cheetah's average speed in miles per hour -------------------------------------- (300 ft/2.92 sec)(1 mi/ 5280 ft)(3600 sec/1 hour) = {300*1*3600 miles)/(2.92*5280*1 hour) = 91080000 miles/15417.6 hours = 70.05 mph ================ Cheers, Uditkulka